Please help!!!

A ball is thrown upward with an initial speed of 50ft/sec. It's height after t seconds is given by h=50t - 16t^2.
A) What is the height of the ball after 1.3 seconds?
B) When does the ball hit the ground?
For A, the answer I worked out is 37.96 feet.
For B, I worked out 25/8 seconds = t, which doesn't seem right. Can anyone help?
Thank you!!!!

Answer by stanbon(75887)

A ball is thrown upward with an initial speed of 50ft/sec. It's height after t seconds is given by h=50t - 16t^2.
A) What is the height of the ball after 1.3 seconds?
For A, the answer I worked out is 37.96 feet.
That is correct.
---------------------------
B) When does the ball hit the ground?
Let h = 0.
0 = 50t -16t^2
Factor to get:
2t(25 - 8t) = 0
t = 0 or t = 25/8 sec.
Comment: that tells you the ball is on the ground at the beginning
and is back on the ground after 25/8 = 3 1/8 seconds.
------------------------
For B, I worked out 25/8 seconds = t, which doesn't seem right. Can anyone help?
You are correct
==========================
Cheers,
Stan H.